OpenStudy (anonymous):

Solve the equation 3x + 1/3 = 2 for x

OpenStudy (anonymous):

3x+1 = 6 3x = 5 x = 5/3

OpenStudy (anonymous):

Simplify the equation to get x on its own. 3x + 1/3 = 2 3x = 2 - 1/3 x = (2 - 1/3)/3

OpenStudy (anonymous):

It's not 5/3.

OpenStudy (smarty):

Multiply both sides by 3 to get rid of the fraction: 9x+1=6 9x=6-1 9x=5 x=5/9

OpenStudy (anonymous):

thanks guys